OpenStudy (anonymous):

Please help me! Medal+fanning! f(x) = 72(1.25)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. (5 points)

OpenStudy (anonymous):

I know it's increasing. I just don't understand by how much?

OpenStudy (anonymous):

@Whitemonsterbunny17 could u help? :)

whitemonsterbunny17 (whitemonsterbunny17):

Oh, I'm not very good at this kind of math, sorry. :/ Maybe @Novalynne can help though. (:

OpenStudy (anonymous):

Okay, thanks so much! =^.^=

whitemonsterbunny17 (whitemonsterbunny17):

No problem!! *^.^*

whitemonsterbunny17 (whitemonsterbunny17):

Bunnies wishes you the best of luck with finding your answer, and with your future studies!!! *^o^* /) /) O( ^.^)o *:・゚✧.゚・*:・゚ C(")(") ☁☁☁☁☁☁☁☁☁☁☁☁☁☁ Also, thank you for using Open Study!!! *^.^*

OpenStudy (anonymous):

Thanks, haha!! :)

OpenStudy (anonymous):

Give me like 30 minutes and if you still need help I'll see if I can figure it out :) I'm in the middle of a timed test.

OpenStudy (anonymous):

Okay, I'll try to wait. :)

OpenStudy (anonymous):

Is it increasing by 75%, I'm confused.

OpenStudy (anonymous):

Okay I finished early! This doesn't look very familiar to me.. what level math class are you in?

OpenStudy (anonymous):

I remember doing functions before but nothing like this.

OpenStudy (anonymous):

I'm in Algebra 1 only because I need segment 1 but they made this course harder. :(

OpenStudy (anonymous):

are u in 7th grade

OpenStudy (anonymous):

Yeah I used functions in Algebra 1 before but not like this. Sorry! :(

OpenStudy (anonymous):

@OrangeMaster Do you know how to do it?

OpenStudy (anonymous):

mia r u in 7th

OpenStudy (anonymous):

No, I'm in 10th. :( It's fine, know anyone who could help? I just need this question and one more.

OpenStudy (anonymous):

No, sorry though. I'm only in 9th grade.

OpenStudy (anonymous):

I keep doing that and nothing comes out right. ;(

OpenStudy (anonymous):

This is 9th grade math, I just didn't have segment 1 done yet.

OpenStudy (anonymous):

@Hero can you pleasseee help?

OpenStudy (anonymous):

Well when you solve it you get f(x)=90.x So I'm assuming you solve for the % difference between 72 and 90?

OpenStudy (anonymous):

I thought it was 75% but I wasn't sure.

OpenStudy (anonymous):

@jlg030597 do u know?

OpenStudy (anonymous):

Well to find the difference you have to first subtract 72 from 90 so 90-72=?

OpenStudy (anonymous):

I'm gonna close this and re ask again :)

OpenStudy (anonymous):

90-72= 18 The next step is the find the average of 90 and 72. To find that you do 90+72/2=117 Now you do 117/18=6.5 So the percent increase would be 6.5%

OpenStudy (anonymous):

I'm not sure If it's right though because I'm still not sure if you're supposed to find the difference between 72 &90.

OpenStudy (anonymous):

Thanks so much! I have to go now! I got a percentage change of 25% and 30%.